# RadarDistill: Boosting Radar-based Object Detection Performance via Knowledge Distillation from LiDAR Features

### 08 Mar 2024 v1

@@ Abstract
The inherent noisy and sparse characteristics of radar data pose challenges in finding effective representations for 3D object detection. In this paper, we propose RadarDistill, a novel knowledge distillation (KD) method, which can improve the representation of radar data by leveraging LiDAR data. RadarDistill successfully transfers desirable characteristics of LiDAR features into radar features using three key components: Cross-Modality Alignment (CMA), Activation-based Feature Distillation (AFD), and Proposal-based Feature Distillation (PFD). CMA enhances the density of radar features through multiple layers of dilation operations, effectively addressing the challenges of inefficient knowledge transfer from LiDAR to radar. AFD is designed to transfer knowledge from significant areas of the LiDAR features, specifically those regions where activation intensity exceeds a predetermined threshold. PFD guides the radar network to mimic LiDAR network features in the object proposals for accurately detected results while moderating features for misdetected proposals like false positives. Our comparative analyses conducted on the nuScenes datasets demonstrate that RadarDistill achieves state-of-the-art (SOTA) performance for radar-only object detection task, recording 20.5% in mAP and 43.7% in NDS. Also, RadarDistill significantly improves the performance of the camera-radar fusion model.
Figure 1: Illustration of RadarDistill. Through knowledge transfer from LiDAR features to radar features, our RadarDistill converts sparse and noisy radar features into dense and semantically rich LiDAR-like features for better object detection.
@@ 1 Introduction
This paper focuses on improving the performance of radar-based 3D object detection using deep neural networks. We note that the limited performance of radar is largely due to the challenges in finding effective representations, given the sparse and noisy nature of radar measurements. Inspired by the remarkable success of deep models that encode LiDAR point clouds, our goal is to transfer the knowledge extracted from a LiDAR-based model to enhance a radar-based model. Recently, knowledge distillation (KD) techniques have shown success in transferring knowledge from one sensor modality to another, thereby refining the representation quality of the target model. To date, various KD methods have been introduced in the literature [8, 4, 11, 3, 41, 16, 32, 40]. Cross-modality knowledge distillation can be broadly divided into two approaches based on the modality of the student model used. When employing a camera as the student model, both depth and shape information are transferred from the teacher model to the student model. BEVDistill [3] transforms both LiDAR and cameras features into a BEV format, enabling the transfer of spatial knowledge from LiDAR features to camera features. DistillBEV [32] utilizes the prediction results from either LiDAR or LiDAR-camera fusion models to distinguish between foreground and background, guiding the student model to focus on KD in essential areas. S2M2-SSD [40] determines key areas based on the student model’s predictions and transfers information obtained from a LiDAR-camera fusion model in the key areas. Apart from these approaches, UniDistill [41] employs a universal cross-modality framework that enables knowledge transfer among diverse modalities. This framework is adaptable to different modality pairings, including camera-to-LiDAR, LiDAR-to-camera, and (camera+LiDAR)-to-camera settings.
@@
In this paper, we propose RadarDistill, a novel KD framework designed to enhance the representation of radar data, leveraging LiDAR data. Our study shows that by employing a radar encoding network as a student network and a LiDAR encoding network as a teacher network, our KD framework effectively produces radar features akin to the dense and semantically rich features derived from LiDAR data for better object detection. Although both LiDAR data and its encoding network are used to enhance radar features during the training phase, they are not required in the inference phase.
@@
Our proposed RadarDistill is designed based on three main ideas: 1) Cross-Modality Alignment (CMA), 2) Activation-based Feature Distillation (AFD), and 3) Proposal-based Feature Distillation (PFD). Our study indicates that transferring knowledge from LiDAR to radar features is difficult due to the inherent sparsity of radar data, which complicates the alignment with the more densely distributed LiDAR features. To address this problem, CMA boosts the student network’s capacity and simultaneously increases the ratio of activated radar features by implementing multiple layers of dilation operations.
@@
The proposed AFD and PFD minimize the distribution gap between the intermediate features produced by the radar and LiDAR encoding networks. Initially, AFD conducts Activation-aware Feature Matching on low-level features. Specifically, it divides both radar and LiDAR features into active and inactive regions according to activation intensity of each features and constructs KD losses for each region separately. By assigning greater importance to the KD loss linked with active regions, AFD enables the network to concentrate on transferring knowledge for significant features.
@@ 2.2 Knowledge Distillation
Figure 2: Overall architecture of RadarDistill. The input point clouds from each modality are initially processed through Pillar Encoding and Sparse 2D CNN to extract low-level features. CMA processes the low-level features of the radar to make them denser and align them with the features of LiDAR. AFD distinguishes between active and inactive regions in both radar and LiDAR features and minimizes their KD losses using different importance factors. Subsequently, proposal-level features are identified using the detection head, and PFD reduces the distribution gap between proposal features obtained from the radar and LiDAR networks. LiDAR data is only needed during the training phase to enhance radar features and not during inference.
@@ 3.1 Preliminary
We used PillarNet [23] as our baseline model for both LiDAR and radar detectors. PillarNet initially generates regular pillars from the point clouds, subsequently producing sparse 2D pillar features denoted as $F^{2D}_{ldr}$ for LiDAR and $F^{2D}_{rdr}$ for radar through the pillar feature network. Then, we define the low-level BEV features denoted by $F^{l}_{ldr}$ and $F^{l}_{rdr}\in\mathbb{R}^{C\times\frac{H}{8}\times\frac{W}{8}}$ obtained by encoding 2D pillar features with a 2D sparse convolution-based encoder (SparseEnc), as follows
$F^{l}_{mod}=SparseEnc(F^{2D}_{mod}),$ (1)
where low-level BEV features are sparse features derived from processing through SPConv [6] and $mod$ represents the detector modality, either ldr or rdr. We describe the high-level BEV features denoted by $F^{h}_{ldr}$ and $F^{h}_{rdr}\in\mathbb{R}^{C\times\frac{H}{8}\times\frac{W}{8}}$, formed by encoding low-level BEV features $F^{l}_{ldr}$ and $F^{l}_{rdr}$ using a 2D dense convolution-based encoder (DenseEnc) as
$F^{h}_{mod}=DenseEnc(F^{l}_{mod}).$ (2)
The high-level BEV features are processed through a centerpoint[38]-based head (CenterHead) to generate the final prediction heatmaps for classification, regression, and IoU scoring, represented as $H^{cls}$, $H^{reg}$ and $H^{IoU}$. The final predictions are obtained from
$H^{cls}_{mod},H^{reg}_{mod},H^{IoU}_{mod}=CenterHead(F^{h}_{mod}).$ (3)
@@ 3.2 Cross-Modality Alignment
Our goal is to reduce the disparity between radar and LiDAR due to the differences in point cloud density via CMA. PillarNet [23] employs SPConv [6] to generate low-level features from only non-empty pillars. Notably, when comparing the number of non-empty pillars, radar’s non-empty pillars ($F^{l}_{rdr}$) comprise only $11\%$ of the total count of LiDAR’s non-empty pillars ($F^{l}_{ldr}$). This significant disparity in the quantity of non-empty pillars necessitates an alignment of the two modalities, especially since the information from LiDAR’s non-empty pillars cannot be directly transferred to the corresponding empty pillars in radar data.
Fig. 3 presents the structure of the CMA. Within the component of CMA, the down block applies deformable convolution [5] to selectively extract essential features during downsampling, and then it aggregates these features using the ConvNeXt V2 blocks [33]. In the up block, dilation is applied using 2D transposed convolution to densify the features in surrounding areas. The aggregation module that includes concatenation followed by a 1$\times$1 convolution layer combines different features and operates similarly to a skip connection.
We downsample the 1/8 resolution features $F^{l}_{rdr}$ to produce $F^{l^{\prime(1)}}_{rdr}$. These $F^{l^{\prime(1)}}_{rdr}$ features are upsampled and combined with $F^{l}_{rdr}$ in the aggregation module, forming the aligned BEV features $F^{a^{(1)}}$. Subsequently, we reduce the resolution of $F^{l^{\prime(1)}}_{rdr}$ to 1/32 and then increase it to 1/16. These upsampled features are integrated with the downsampled $F^{a^{(1)}}$, to produce $F^{l^{\prime(2)}}_{rdr}$. Finally, we upsample the $F^{l^{\prime(2)}}_{rdr}$ and fuse it with $F^{a^{(1)}}$ in the aggregation module, which generates the second aligned BEV features $F^{a^{(2)}}$. In the CMA process, we acquire two sets of aligned features at 1/8 resolution. We apply multiple AFD to these aligned features to enhance their similarity to LiDAR’s low-level features. Consequently, as observed in Fig. 1, the radar’s low-level features become denser after being processed through the CMA. A comprehensive description of the CMA’s component is available in the Supplementary Materials.
@@ 3.3 Activation-based Feature Distillation
AFD strategically aligns the low-level features of radar and LiDAR through Activation-aware Feature Matching. This process matches the activation patterns of radar features to those of LiDAR, thus bridging the gap in their distribution of active features. Additionally, AFD is applied iteratively to each pair of the $n$-th aligned features, $F^{a^{(n)}}$ and $F^{l}_{ldr}$, ensuring a more precise and effective alignment between the radar and LiDAR features. First, for both LiDAR and radar features, we define “active regions” as areas where each radar and LiDAR feature intensity values, summed in the channel direction, are greater than 0. Conversely, areas where these sums are less than 0 are classified as “inactive regions”. We define active mask $M^{mod}_{i,j}$ to present these regions and obtain the active masks $M^{a^{(n)}}_{i,j}$and $M^{ldr}_{i,j}$ for $F^{a^{(n)}}$and $F^{l}_{ldr}$ correspondingly. These masks are obtained from
$F^{mod}_{i,j}=\sum^{C}_{c=1}F^{mod}_{c,i,j},$ (4)
$M^{mod}_{i,j}=\begin{cases}1,&if\,(F^{mod}_{i,j}>0),\\ 0,&otherwise,\end{cases}$ (5)
where $F\in\mathbb{R}^{C\times\frac{H}{8}\times\frac{W}{8}}$ represents BEV features, the indices i,j,c correspond to height, width, and channel, and $mod$ indicates the modality of features, comprising either ldr or a${}^{(n)}$. The distillation region is adaptively separated into two categories: the active regions (AR) where both radar and LiDAR are active and the inactive regions (IR) where radar is active while LiDAR is inactive. There is an imbalance between the AR and IR due to the sparse and noisy characteristics of radar. This can disturb training by focusing on imitating the dominant IR. Therefore, we apply a relative adaptive weight based on the number of pixels in each region. AR is assigned a weight of 1 as it is a smaller yet crucial region while IR is given a relative adaptive weight denoted by $\rho$ as determined by the ratio of pixel counts between AR and IR. We define a separate mask $M^{sep^{(n)}}$ for distinguishing the regions, specifically between $F^{a^{(n)}}$and $F^{l}_{ldr}$ as
$AR^{(n)}$ $=\,(M^{a^{(n)}}=1)\,\&\,(M^{ldr}=1),$ (6)
$IR^{(n)}$ $=\,(M^{a^{(n)}}=1)\,\&\,(M^{ldr}=0),$ (7)
$M^{sep^{(n)}}_{i,j}$ =$=\begin{cases}\alpha,&if\,(i,j)\in AR^{(n)},\\ \rho^{(n)}\times\beta,&if\,(i,j)\in IR^{(n)},\\ 0,&\,otherwise,\end{cases}$ (8)
$\rho^{(n)}$ $=\frac{N_{AR^{(n)}}}{N_{IR^{(n)}}},$ (9)
where $\rho^{(n)}$ represents the relative importance of the $IR^{(n)}$ in comparison to $AR^{(n)}$, $\alpha$ and $\beta$ are the balancing parameters between the separated regions, and $N$ is the number of pixels. In this way, the distillation region and the $\rho^{(n)}$ value are adaptively adjusted for each training, and distillation is performed with different weights for AR and IR.
Figure 3: Detailed structure of the proposed CMA module.
After decomposing the distillation region with AR and IR, the $n$-th aligned features $F^{a^{(n)}}_{rdr}$ are trained to imitate $F^{l}_{ldr}$. The distillation loss $L^{(n)}_{low}$ for the $n$-th activation features is given by
$L^{(n)}_{low}$ $=\sum^{C}_{c=1}\sum^{H}_{i=1}\sum^{W}_{j=1}M^{sep^{(n)}}_{i,j}(F^{ldr}_{c,i,j}-F^{a^{(n)}}_{c,i,j})^{2},$ (10)
where $F^{ldr}$ denotes the LiDAR’s low-level BEV features $F^{l}_{ldr}$.
The active mask loss is applied between the predicted mask, obtained by applying a sigmoid function to $F^{a(n)}$, and $M^{ldr}$ to ensure the radar’s active mask closely mimics LiDAR. The active mask loss $L^{(n)}_{mask}$ is calculated as
$L^{(n)}_{mask}$ $=\sum^{H}_{i=1}\sum^{W}_{j=1}|M^{ldr}_{i,j}-\sigma(F^{a^{(n)}}_{i,j})|,$ (11)
where $\sigma(\cdot)$ denotes a sigmoid function.
We apply AFD twice to the aligned BEV features to more accurately imitate $F^{l}_{ldr}$, and use the average of losses calculated from these multiple AFD as the final AFD loss. The final AFD loss function is given by
$L_{AFD}$ $=\frac{1}{2}\sum^{2}_{n=1}(L^{(n)}_{low}+L^{(n)}_{mask}).$ (12)

### 05 Apr 2024 v2

@@ Abstract
The inherent noisy and sparse characteristics of radar data pose challenges in finding effective representations for 3D object detection. In this paper, we propose RadarDistill, a novel knowledge distillation (KD) method, which can improve the representation of radar data by leveraging LiDAR data. RadarDistill successfully transfers desirable characteristics of LiDAR features into radar features using three key components: Cross-Modality Alignment (CMA), Activation-based Feature Distillation (AFD), and Proposal-based Feature Distillation (PFD). CMA enhances the density of radar features by employing multiple layers of dilation operations, effectively addressing the challenge of inefficient knowledge transfer from LiDAR to radar. AFD selectively transfers knowledge based on regions of the LiDAR features, with a specific focus on areas where activation intensity exceeds a predefined threshold. PFD similarly guides the radar network to selectively mimic features from the LiDAR network within the object proposals. Our comparative analyses conducted on the nuScenes datasets demonstrate that RadarDistill achieves state-of-the-art (SOTA) performance for radar-only object detection task, recording 20.5% in mAP and 43.7% in NDS. Also, RadarDistill significantly improves the performance of the camera-radar fusion model.
Figure 1: Illustration of Proposed RadarDistill. Our RadarDistill method facilitates knowledge transfer from LiDAR features to radar features, enhancing the quality of radar features for Bird’s Eye View (BEV) object detection.
@@ 1 Introduction
This paper focuses on improving the performance of radar-based 3D object detection using deep neural networks. We note that the limited performance of radar is largely due to the challenges in finding effective representations, given the sparse and noisy nature of radar measurements. Inspired by the remarkable success of deep models that encode LiDAR point clouds, our goal is to transfer the knowledge extracted from a LiDAR-based model to enhance a radar-based model.
@@
Recently, knowledge distillation (KD) techniques have shown success in transferring knowledge from one sensor modality to another, thereby improving the representation quality of the target model. To date, various KD methods have been introduced in the literature [8, 4, 11, 3, 42, 16, 33, 41]. BEVDistill [3] transforms both LiDAR and cameras features into a BEV format, enabling the transfer of spatial knowledge from LiDAR features to camera features. DistillBEV [33] utilizes the prediction results from either LiDAR or LiDAR-camera fusion models to distinguish between foreground and background, guiding the student model to focus on KD in essential areas. S2M2-SSD [41] determines key areas based on the student model’s predictions and transfers information obtained from a LiDAR-camera fusion model in the key areas. Apart from these approaches, UniDistill [42] employs a universal cross-modality framework that enables knowledge transfer among diverse modalities. This framework is adaptable to different modality pairings, including camera-to-LiDAR, LiDAR-to-camera, and (camera+LiDAR)-to-camera settings.
@@
In this paper, we propose RadarDistill, a novel KD framework designed to enhance the representation of radar data, leveraging LiDAR data. Our study shows that by employing a radar encoding network as a student network and a LiDAR encoding network as a teacher network, our KD framework effectively produces radar features akin to the dense and semantically rich LiDAR features. Although both LiDAR data and its encoding network are used to enhance radar features during the training phase, they are not required in the inference phase.
@@
Our proposed RadarDistill is designed based on three main ideas: 1) Cross-Modality Alignment (CMA), 2) Activation-based Feature Distillation (AFD), and 3) Proposal-based Feature Distillation (PFD). Our study indicates that transferring knowledge from LiDAR to radar features is difficult due to the inherent sparsity of radar data, which complicates the alignment with the more densely distributed LiDAR features. To address this problem, CMA boosts the student network’s capacity and simultaneously increases the ratio of active radar features by implementing multiple layers of dilation operations.
@@
The proposed AFD and PFD minimize the distribution gap between the intermediate features produced by the radar and LiDAR encoding networks. Initially, AFD conducts Activation-aware Feature Matching on low-level features. Specifically, it divides both radar and LiDAR features into active and inactive regions according to activation intensity of each features and constructs KD losses for each region separately. By assigning greater importance to the KD loss linked with sparsely distributed active regions, AFD enables the network to concentrate on transferring knowledge for significant features.
@@
Figure 2: Overall architecture of RadarDistill. The input point clouds from each modality are independently processed through Pillar Encoding followed by SparseEnc to extract low-level BEV features. CMA is then employed to densify the low-level BEV features in the radar branch. AFD then identifies active and inactive regions based on both radar and LiDAR features and minimizes their associated distillation losses. Subsequently, PFD conducts knowledge distillation based on proposal-level features obtained from DenseEnc. Note that the LiDAR branch is solely utilized during the training phase to enhance the radar pipeline and is not required during inference.
@@ 3.1 Preliminary
We use PillarNet [23] as our baseline model for both LiDAR and radar detectors. PillarNet organizes both radar and LiDAR point clouds using a 2D pillar structure of the same size. It then generates two BEV pillar features $F^{2D}_{ldr}$ and $F^{2D}_{rdr}$, utilizing separate pillar encoding networks in the LiDAR and radar branches, respectively. We define the low-level BEV features $F^{(l)}_{ldr}$ and $F^{(l)}_{rdr}\in\mathbb{R}^{C\times\frac{H}{8}\times\frac{W}{8}}$ obtained by encoding the pillar features $F^{2D}_{ldr}$ and $F^{2D}_{rdr}$ as
$F^{(l)}_{mod}=SparseEnc(F^{2D}_{mod}),$ (1)
where $SparseEnc(\cdot)$ denotes a 2D sparse convolution-based encoder (SparseEnc) [6] and $mod$ represents the detector modality, i.e., $mod\in\{ldr,rdr\}$. Similarly, we describe the high-level BEV features $F^{(h)}_{ldr}$ and $F^{(h)}_{rdr}\in\mathbb{R}^{C\times\frac{H}{8}\times\frac{W}{8}}$ formed by encoding the low-level BEV features $F^{(l)}_{ldr}$ and $F^{(l)}_{rdr}$, respectively as
$F^{(h)}_{mod}=DenseEnc(F^{(l)}_{mod}),$ (2)
where $DenseEnc(\cdot)$ denotes a 2D dense convolution-based encoder (DenseEnc). The high-level BEV features are further processed through a CenterHead network [39] to generate the final prediction heatmaps for classification, regression, and IoU scoring
$H^{cls}_{mod},H^{reg}_{mod},H^{IoU}_{mod}=CenterHead(F^{(h)}_{mod}).$ (3)
Note that the pipeline in the LiDAR branch serves as the teacher model, while the one in the radar branch acts as the student model in our KD framework. These networks, serving as teacher and student, respectively, do not share weights.
@@ 3.2 Cross-Modality Alignment
The objective of CMA is to enhance the density of radar BEV features, thereby facilitating the transfer of knowledge from LiDAR features to radar features more effectively. We note that the average number of non-empty pillars in radar features $F^{(l)}_{rdr}$ is only approximately 10% of the average number found in LiDAR features $F^{(l)}_{ldr}$. This considerable difference in the number of non-empty pillars poses challenges in aligning features between the two modalities. Specifically, transferring information from non-empty pillars in LiDAR data to corresponding empty pillars in radar data proves to be impractical. To tackle this challenge, CMA is employed to densify radar features. Our empirical study demonstrates that this densification process significantly contributes to the successful knowledge distillation from LiDAR to radar.
Fig. 3 illustrates the architecture of CMA. CMA comprises the Down Block, Up Block, and Aggregation Module. Down Block conducts down-sampling via deformable convolution [5], followed by the ConvNeXt V2 blocks [34]. The Up Block conducts up-sampling operations using a 2D transposed convolution. The Aggregation Module concatenates two input features and applies a 1$\times$1 convolution layer. After applying the Down Block and Up Block operations twice each, in addition to incorporating a side pathway through the Aggregation Module, CMA generates two low-level BEV features, $F^{(l_{1})}_{rdr}$ and $F^{(l_{2})}_{rdr}$, at a $1/8$ resolution. As a result of CMA’s dilation operation, these low-level features exhibit increased density compared to the input radar features. (Refer to Fig. 2 for a visual comparison of the features before and after CMA.) These intermediate features are subsequently utilized for knowledge distillation in the following AFD stage. For a detailed structure of the CMA, refer to the Supplementary Materials.
Figure 3: Detailed structure of the proposed CMA module.
@@ 3.3 Activation-based Feature Distillation
AFD conducts Activation-aware Feature Matching on the low-level features obtained from both radar and LiDAR branches. AFD performs knowledge distillation from the LiDAR BEV features $F^{(l)}_{ldr}$ to the first radar BEV features $F^{(l_{1})}_{rdr}$. Simultaneously, it also performs knowledge distillation from $F^{(l)}_{ldr}$ to the second radar BEV features $F^{(l_{2})}_{rdr}$.
AFD conducts selective knowledge distillation based on regions using active masks. We determine the active masks $M^{(l_{1})}_{rdr}$, $M^{(l_{2})}_{rdr}$, and $M^{(l)}_{ldr}$ using the corresponding densified radar features $F^{(l_{1})}_{rdr}$, $F^{(l_{2})}_{rdr}$, and the low-level LiDAR features $F^{(l)}_{ldr}$, respectively, as
$F^{(l^{\prime})}_{mod,i,j}=\sum^{C}_{c=1}F^{(l^{\prime})}_{mod,c,i,j},$ (4)
$M^{(l^{\prime})}_{mod,i,j}=\begin{cases}1,&{\rm if}\,F^{(l^{\prime})}_{mod,i,j}>0,\\ 0,&{\rm otherwise},\end{cases}$ (5)
where $F\in\mathbb{R}^{C\times\frac{H}{8}\times\frac{W}{8}}$ represents BEV features, $i,j,$ and $c$ correspond to height, width, and channel indices, respectively, $mod\in\{ldr,rdr\}$, and $l^{\prime}\in\{l_{1},l_{2},l\}$.
Using the active masks obtained from (4) and (5), we can identify the active regions (AR) where both radar and LiDAR features are active, and the inactive regions (IR) where radar features are active and LiDAR features are inactive, i.e.,
$AR^{(l_{n})}$ $=\,(M^{(l_{n})}_{rdr}=1)\,\&\,(M^{(l)}_{ldr}=1),$ (6)
$IR^{(l_{n})}$ $=\,(M^{(l_{n})}_{rdr}=1)\,\&\,(M^{(l)}_{ldr}=0),$ (7)
where $n\in[1,2]$. AFD minimizes the difference between radar and LiDAR features in both the AR and IR. AFD is specifically trained to transform radar features to resemble LiDAR features within AR, while simultaneously learning to suppress radar features within IR.
Due to the sparsity of LiDAR data, the AR area is typically much smaller than the IR area. To prevent the model from overly focusing on knowledge distillation within IR, we scale the loss terms associated with AR and IR based on their relative area ratios. We devise adaptive loss weights $W^{(l_{n})}_{sep}$ to balance the loss terms within AR and IR according to their respective pixel counts, i.e.,
$W^{(l_{n})}_{sep,i,j}$ =$=\begin{cases}\alpha,&if\,(i,j)\in AR^{(l_{n})},\\ \rho^{(l_{n})}\times\beta,&if\,(i,j)\in IR^{(l_{n})},\\ 0,&\,otherwise,\end{cases}$ (8)
where $\rho^{(l_{n})}=\frac{N_{AR^{(l_{n})}}}{N_{IR^{(l_{n})}}}$ represents the relative importance of the $IR^{(l_{n})}$ over $AR^{(l_{n})}$, $\alpha$ and $\beta$ are the intrinsic balancing parameters, and $N_{AR^{(l_{n})}}$ and $N_{IR^{(l_{n})}}$ are the number of pixels in $AR^{(l_{n})}$ and $IR^{(l_{n})}$, respectively. The weight associated with IR is determined proportionally by $\rho$, representing the ratio of pixel count for AR to that for IR.
The distillation loss $L^{(n)}_{low}$ for the $n$-th activation features is given by
$L^{(n)}_{low}$ $=\sum^{C}_{c=1}\sum^{H}_{i=1}\sum^{W}_{j=1}W^{(l_{n})}_{sep,i,j}(F^{(l)}_{ldr,c,i,j}-F^{(l_{n})}_{rdr,c,i,j})^{2}.$ (9)
The final AFD loss $L_{AFD}$ is obtained by averaging the distillation loss functions associated with two densified radar features $F^{(l_{1})}_{rdr}$ and $F^{(l_{2})}_{rdr}$
$L_{AFD}$ $=\frac{1}{2}\sum^{2}_{n=1}L^{(n)}_{low}.$ (10)